# Clustering of timed sequences – Application to the analysis of care pathways

### v1

@@ Abstract
Improving the future of healthcare starts by better understanding the current actual practices in hospitals. This motivates the objective of discovering typical care pathways from patient data. Revealing homogeneous groups of care pathways can be achieved through clustering. The difficulty in clustering care pathways, represented by sequences of timestamped events, lies in defining a semantically appropriate metric and clustering algorithms.
In this article, we adapt two methods developed for time series to time sequences: the drop-DTW metric and the DBA approach for the construction of averaged time sequences. These methods are then applied in clustering algorithms to propose original and sound clustering algorithms for timed sequences. This approach is experimented with and evaluated on synthetic and real use cases.
@@ 1 Introduction
Timed sequences, composed of dated events, are types of data frequently encountered in various application domains such as computer log analysis, life course analysis in social science [6], tourism trajectories [1] or healthcare pathway analysis [16]. A peculiarity of this type of data is that they do not have a natural vector representation. In particular, they are formed by two very different dimensions: symbolic and temporal; and, on the other hand, their size generally varies from one example to another (e.g., in a healthcare pathway, some patients have more care events than others). This difficulty in modeling data in a vector space makes it challenging to adapt standard data analysis methods (classification, clustering, or forecasting).
@@
The adaptation of data analysis methods raises the question of formalizing a topology of the objects to be analyzed. Most of the data analytic methods make the assumption that objects lies in a metric space. This topology assumption provides intuitive ways to generalize sparse and noisy observations. Formalizing such a space for timed sequences involves defining a metric between timed sequences. This problem is very similar to that encountered in time series analysis (temporal data where the values are numerical, not event-based, and generally with a regular and dense temporal sampling). Thus, for time series analysis, a large collection of metrics has been proposed and allows them to be analyzed [11, 21]. Then, the challenges becomes to identify the metric that has the properties appropriate for the analysis task one wishes to carry out (e.g., the properties of a distance) and whose semantics will give the expected results. Intuitively, a similarity metric (resp. dissimilarity metric) should produce a high similarity value (resp. low dissimilarity value) when two examples are considered similar by domain experts.
@@
For sequential data, few studies have focused on adapting data analysis methods. [8] and [9] are based on tensor representation of sequential data, while [22] or [12] use embedding representations. But only a few methods represent metric time [15, 6]. Beside these end-to-end methods, some works were dedicated to metrics between temporal sequences [7, 19, 4] or averaging [3, 2, 20]. Despite these proposals, metrics whose semantics can better take into account the temporal dimension are still lacking (and in particular the delays between the occurrence of events). Indeed, most current solutions generally only take into account the sequentiality in the data.
@@
In this work, we extend the drop-DTW metric [5] which was proposed to evaluate distances between sequences by taking into account both sequential aspects, by adapting Dynamic Time Warping [18], and by adding support for meaningless events, which can be dropped. This metric has the semantic characteristics that seem appropriate for our application case – i.e., the clustering of care pathways, but as this metric is not a distance, it does not allow for the construction of an average sequence nor to be used in classical clustering methods such as the K-Means algorithm or hierarchical clustering.
@@
The main contribution of this article is the proposal of an algorithm to compute a mean timed sequences based on drop-DTW. This algorithm is inspired by the DBA method [14] for time series. The construction of an average sequence meets both the needs of clustering algorithms (hierarchical clustering or K-means) and the needs of interpretation of sequence clusters.
@@
This method is applied to discover typical care-pathways from Electronic Health Record (EHR) datasets. In the context of the OPTISOINS project [^1] https://www.bernoulli-lab.fr/project/optisoins/, we were particularly interested in patients who had undergone pulmonary resection surgery for oncological reasons. Care pathways were reconstructed from data collected in the EHR data warehouse of the Greater Paris Hospital (AP-HP). The care pathway is the timed sequence of the key stages of managing the patient surgery from diagnosis to post-surgery treatments.
@@ 2 Times Sequences and Probabilistic Timed Sequences
Let $\Sigma$ be a finite set of $n$ event types. A timed sequence is a sequence of pairs $S=\langle(s_{i},t_{i})\rangle$ where $s_{i}\in\Sigma$ is an event type and $t_{i}\in\mathbb{R}$ is the event’s timestamp. $S[i]=(s_{i},t_{i})$ denotes the $i$-th event in the sequence. $|S|$ is the length of the timed sequences, i.e. its number of events. By convention, events in a sequence are ordered by their timestamps. Several events may occur at the same time if their event types are different, i.e. $(s_{i},t)=(s_{j},t)\Rightarrow s_{i}=s_{j}$.
We propose embedding the space of timed sequences, introduced in the previous section, into a larger space where events are represented by their one-hot encoding, which is a vector representation. Instead of representing a sequence of events, we represent a sequence of event types distributions:
@@
Figure 1: Illustration of the timed sequences of Example 1 (on the left) and the cost matrix for the alignment of $s_{1}$ and $s_{2}$ according to the drop-DTW (on the right). The alignment is illustrated by the grey bars on the left and by the colored cells in the matrix (on the right). The red-crosses illustrate that $D$ has been dropped in this alignment, the cells at the corner are dark because they to no satisfy the constraint $\tau=3.5$.
The example 1 illustrates the representation of a timed sequence as a probabilistic timed sequence. Conversely, most of the probabilistic timed sequences can not be represented by a time sequence.
@@
Let $\Sigma={A,B,C,D}$ be a set of event types, and $s_{1}=\langle(A,1),(C,2),(B,4.5)\rangle$, $s_{2}=\langle(C,0),(C,2),(D,3),(A,4),(B,5)\rangle$, and $s_{3}=\langle(A,0),(C,1),(C,2),(B,4)\rangle$ be three timed sequences illustrated in Figure 1. These sequences can be represented as probabilistic sequences:
@@ 3.1 DTW and drop-DTW
Dynamic Time Warping (DTW) [18] computes the optimal alignment between two series subject to certain constraints. Here, we adapt these definitions to the case of timed sequences. Let $X=\langle(x_{1},t^{x}_{1}),\dots,(x_{N},t^{x}_{N})\rangle$ and $Z=\langle(z_{1},t^{z}_{1})\dots,(z_{K},t^{z}_{K})\rangle$ be the input sequences, where $N,K$ are the respective lengths of the sequences. An alignment between the two sequences is defined as a binary matrix, $M\in\{0,1\}^{K\times N}$ where $M_{i,j}=1$ if $z_{i}$ is matched to $x_{j}$, and 00 otherwise. Matching an element $z_{i}$ to an element $x_{j}$ incurs a cost $C_{i,j}$ between the $i$-th and $j$-th elements of the sequences. DTW finds the alignment $M^{*}$ between sequences $Z$ and $X$ that minimizes the global cost:
$M^{*}=\operatorname*{argmin}_{M\in\mathcal{M}}\langle M,C\rangle=\operatorname*{argmin}_{M\in\mathcal{M}}\sum_{i,j}M_{i,j}C_{i,j},$
where $\langle M,C\rangle$ is the Frobenius inner product and $\mathcal{M}$ is the set of all feasible alignments satisfying the following constraints: monotonicity, continuity, and end-point correspondence ($M_{1,1}=M_{K,N}=1,\forall M\in\mathcal{M}$).
drop-DTW [5] extends the set of feasible alignments, $\mathcal{M}$, to those adhering only to the monotonicity constraint. Consequently, unlike DTW, events can be dropped from the alignment process. drop-DTW finds the optimal temporal alignment while allowing for the removal of certain elements. To account for unmatched elements in the alignment cost calculation, Dvornik et al. [5] introduce a new drop cost: $\delta\in\mathbb{R}^{+}$. [^2] In the original approach of Dvornik et al., the cost value depends on the event type. We use a constant value for the sake of simplification. The optimal matching can then be defined as follows:
$M^{*}=\operatorname*{argmin}_{M\in\overline{\mathcal{M}}}\langle M,C\rangle+\delta.\left(P_{z}(M)+P_{x}(M)\right)$
where $\overline{\mathcal{M}}$ is the set of binary matrices satisfying only the monotonicity constraint, and $P_{x}(M)$ is a vector whose $j$th element is $1$ if $M_{:,j}=\bm{0}$ and 00 otherwise; $P_{z}(M)$ is defined similarly, but along the rows. Similar to DTW, drop-DTW can be efficiently evaluated using dynamic programming techniques.
@@ 3.2 Adaptation of drop-DTW for timed sequences
The proposed adaptation of drop-DTW focuses on proposing a metric between events and adding temporal constraints to the drop-DTW algorithm. The Dynamic Time Warping algorithm requires defining a cost between two events to compare sequences. To take into account both time and events, we decided to weigh: a time distance cost $c_{t}$, and an event difference cost $c_{e}$. Having introduced the probabilistic representation of timed sequences, the distance between two events $e_{a}=(\bm{d}_{a}=[d^{1}_{a},\dots d^{n}_{a}],t_{a})$ and $e_{b}=(\bm{d}_{b}=[d^{1}_{b},\dots d^{n}_{b}],t_{b})$ is given by:
@@
where $p_{t},p_{e}\in\mathbb{R}_{+}$ are weights of two terms: a distance between probabilistic events and a time gap. The choice of this metric turns out to consider that time events are in an Euclidean space of dimension $n+1$.
We can now adapt the metric between two timed sequences. We have chosen to use drop-DTW for its ability to handle aberrant or missing events. Algorithm 1 presents the method for calculating the proposed distance between two timed sequences, where the cost matrix is computed from the event distance, $d$. The algorithm finds the best alignment of sequence events to minimize the overall metric. The use of dynamic programming technique ensures the efficiency of the method. For the sake of computing efficiency, we also add constraints on the possible event pairings by allowing matches only between events that are close in time, i.e. with a delay between their occurrence lower than a predefined threshold $\tau$.
@@
1 $D^{+}_{0,0}\leftarrow 0$ ; $D^{+}_{i,0}\leftarrow+\infty$ ; $D^{+}_{0,j}\leftarrow+\infty$ ;
2 $D^{-}_{0,0}\leftarrow 0$ ; $D^{-}_{i,0}\leftarrow i\times\delta$ ; $D^{-}_{0,j}\leftarrow j\times\delta$ ;
@@ 8                   $D_{i,j}\leftarrow+\infty$;
Considering the sequences from Example 1, sequences $s_{1}$ and $s_{2}$ seem close and may be of interest to be grouped together, while sequence 3 seems very different. Without drop-cost, sequences 2 and 3 would be grouped before sequences $s_{1}$ and $s_{2}$. By using the drop-cost, the last event of sequence 3 would not be considered, and the cost would simply be replaced by $\delta$. Thus, they would be judged more similar than sequences 2 and 3.
@@ 4 Construction of an Average Timed Sequence
Computing an average time d sequence consistent with the notion of similarity is essential for clustering algorithms. In the following, we propose a method inspired by the DBA algorithm [13] to create average sequences in terms of drop-DTW, taking into account the possibility of removal. DBA addresses the problem of generating an average time series in terms of DTW – which does not hold distance properties. In this case, the generalized average is defined as a time series that minimizes the global inertia (see Equation 1).
DBA iteratively refines a (randomly initialized) series to minimize its quadratic distance (DTW) with respect to the series to be averaged. At each iteration, the algorithm performs “vertical” centroids of the aligned elements of the time series. The definition of a “vertical” centroid corresponds to averaging series real values. Our approach enables the adaptation of timed sequences with event type distributions.

### v2

@@ Abstract
Improving the future of healthcare starts by better understanding the current actual practices in hospital settings. This motivates the objective of discovering typical care pathways from patient data. Revealing typical care pathways can be achieved through clustering. The difficulty in clustering care pathways, represented by sequences of timestamped events, lies in defining a semantically appropriate metric and clustering algorithms.
In this article, we adapt two methods developed for time series to the clustering of timed sequences: the drop-DTW metric and the DBA approach for the construction of averaged time sequences. These methods are then applied in clustering algorithms to propose original and sound clustering algorithms for timed sequences. This approach is experimented with and evaluated on synthetic and real-world data.
keywords:
metric, temporal data, sequential data, medical informatics
journal: Knowledge and Data Engineering
@@ 1 Introduction
Timed sequences, composed of dated events, are types of data frequently encountered in various application domains such as computer log analysis, life course analysis in social science (Gabadinho et al., 2011), tourism trajectories (Bernard et al., 2018) or healthcare pathway analysis (Rama et al., 2019). A specificity of this type of data is that they have no natural vector representation. In particular, they are formed by two very different dimensions: symbolic and temporal; and, on the other hand, their size generally varies from one example to another (e.g., in a healthcare pathway, some patients have more care events than others). This difficulty in modeling data in a vector space makes it challenging to adapt standard data analysis methods (classification, clustering, or forecasting).
@@
The adaptation of data analysis methods for timed sequences raises the question of formalizing a topology of the objects to be analyzed. Most of the data analytic methods make the assumption that objects lie in a metric space. This topology assumption provides intuitive ways to generalize sparse and noisy observations. Formalizing such a space for timed sequences involves defining a metric between timed sequences. This problem is very similar to one encountered in time series analysis (temporal data where values are numerical, not event-based, and generally with a regular and dense temporal sampling). For time series analysis, a large collection of metrics has been proposed and allows them to be analyzed (Holder et al., 2024; Tavenard et al., 2020). The challenges turns out to identify a metric that has the properties appropriate for the analysis task one wishes to carry out (e.g., the properties of a distance) and whose semantics will give the expected results. Intuitively, a similarity metric (resp. dissimilarity metric) should produce a high similarity value (resp. low dissimilarity value) when two examples are considered similar by domain experts.
@@
For sequential data, few studies have focused on adapting data analysis methods. Gay et al. (2015) and H. Sebia (2024) are based on tensor representation of sequential data, while Yu et al. (2020) or Nguyen et al. (2018) use embedding representations. Nonetheless, only a few methods handle event logs with timestamped events dates (Rama et al., 2019; Gabadinho et al., 2011). The TraMineR (Gabadinho et al., 2011) software has been frequently used for epidemiological studies, for instance by Le Meur et al. (2019). It enables the user to cluster or classify event-based temporal data with different metrics. Nonetheless, this method represents an event-based data as a state-sequence, i.e., a longitudinal sequence which associate at each instant a unique state. One advantage of this tool is to propose an end-to-end analysis of the data. Another end-to-end method for timed sequences is AliClu (Rama et al., 2019) which combine the definition of a time warping metric with a hierarchical clustering algorithm. Contrary to TraMineR, it does not allow the change of the metric. Beside these end-to-end methods, some works were dedicated to metrics between temporal sequences (Garreau et al., 2014; Schneider et al., 2022; Compagnon et al., 2021) or averaging (Chen et al., 2017; Besnard and Guyet, 2023; Tatti and Vreeken, 2012). Despite these proposals, metrics whose semantics can better take into account the temporal dimension are still lacking. In particular, the delays between the occurrence of events is not handle. Most current solutions generally only take into account the sequentiality in the data.
@@
In this work, we extend the drop-DTW metric (Dvornik et al., 2021) which was proposed to evaluate distances between sequences by taking into account both sequential aspects, by adapting Dynamic Time Warping (Sakoe and Chiba, 1978), and by adding support for meaningless events, which can be dropped. This metric has the semantics that seem appropriate for our application case – i.e., the clustering of care pathways, but as this metric is not a distance, it does not allow for the construction of an average sequence nor to be used in classical clustering methods such as the K-Means algorithm or hierarchical clustering.
@@
The main contribution of this article is the proposal of an algorithm to compute a mean timed sequence based on drop-DTW. This algorithm is inspired by the DBA method (Petitjean et al., 2011) for time series. The construction of an average sequence meets both the needs of clustering algorithms (hierarchical clustering or K-means) and the needs of interpretation of sequence clusters (definition of a representative).
@@
This method is applied on a Electronic Health Records (EHR) dataset in order to discover typical care-pathways. In the context of the OPTISOINS project [^1] https://www.bernoulli-lab.fr/project/optisoins/, we were particularly interested in patients who had undergone pulmonary resection surgery for oncological reasons. Care trajectories were reconstructed from data collected in the EHR data warehouse of the Greater Paris Hospital (AP-HP). The care pathway is a timed sequence representing the key stages of managing the patient surgery from diagnosis to post-surgery treatments.
@@ 2 Times Sequences and Probabilistic Timed Sequences
Let $\Sigma$ be a finite set of $n$ event types. A timed sequence is a sequence of pairs $S=\langle(s_{i},t_{i})\rangle$ where $s_{i}\in\Sigma$ is an event type and $t_{i}\in\mathbb{R}$ is the event’s timestamp. $S[i]=(s_{i},t_{i})$ denotes the $i$-th event in the sequence. $|S|$ is the length of the timed sequences, i.e. its number of events. By convention, events in a sequence are ordered by their timestamps. Several events may occur at the same time only if their event types are different, i.e. $(s_{i},t)=(s_{j},t)\Rightarrow s_{i}=s_{j}$.
We propose embedding the space of timed sequences into a larger space where events are represented by their one-hot encoding, which is a vector representation. Instead of representing a sequence of events, we represent a sequence of event types distributions:
@@
Figure 1: Illustration of the timed sequences of Example 1 (on the left) and the cost matrix for the alignment between $s_{1}$ and $s_{2}$ according to the drop-DTW (on the right). The alignment is illustrated by the grey bars on the left and by the colored cells in the matrix (on the right). The red-crosses illustrate that $P$ and $S$ has been dropped in this alignment, the cells at the corner are dark because they to no satisfy the constraint $\tau=3.5$. Clinical pathway interpretation: $S$ stands for surgery, $C$ consultation, $P$ physiotherapist session and $R$ radiotherapy session.
Example 1 illustrates the representation of a timed sequence as a probabilistic timed sequence. Conversely, most of the probabilistic timed sequences can not be represented by a time sequence.
@@
Let $\Sigma=\{S,R,C,P\}$ be an ordered set of event types. In the context of clinical care pathways, $S$ may represent a surgery event, $C$ a consultation, $P$ a physiotherapist session, and $R$ a radiotherapy session. We define three timed sequences: $s_{1}=\langle(S,1),(C,2),(R,4.5)\rangle$, $s_{2}=\langle(C,0),(C,2),(P,3),(S,4),(R,5)\rangle$, and $s_{3}=\langle(S,0),(C,1),(C,2),(R,4)\rangle$ be three synthetic timed sequences illustrated in Figure 1. These sequences can be represented as probabilistic sequences:
@@ 3.1 DTW and drop-DTW
Dynamic Time Warping (DTW) (Sakoe and Chiba, 1978) computes the optimal alignment between two series subject to certain constraints. Here, we adapt these definitions to the case of timed sequences. Let $X=\langle(x_{1},t^{x}_{1}),\dots,(x_{N},t^{x}_{N})\rangle$ and $Z=\langle(z_{1},t^{z}_{1})\dots,(z_{K},t^{z}_{K})\rangle$ be two input sequences, where $N,K$ are their respective lengths. An alignment between two sequences is defined as a binary matrix, $\pi\in\{0,1\}^{K\times N}$ where $\pi_{i,j}=1$ if $z_{i}$ is matched to $x_{j}$, and 00 otherwise. Matching an element $z_{i}$ to an element $x_{j}$ incurs a cost $C_{i,j}$ between the $i$-th and $j$-th elements of the sequences. The cost values are user-defined. DTW finds the alignment $\pi^{*}$ between sequences $Z$ and $X$ that minimizes the global cost:
$\pi^{*}=\operatorname*{argmin}_{\pi\in\mathcal{P}}\,\langle\pi,C\rangle=\operatorname*{argmin}_{\pi\in\mathcal{P}}\sum_{i\in[K],\,j\in[N]}\pi_{i,j}C_{i,j},$
where $\langle\pi,C\rangle$ is the Frobenius inner product and $\mathcal{P}$ is the set of all feasible alignments satisfying the following constraints: monotonicity, continuity, and end-point correspondence ($\pi_{1,1}=\pi_{K,N}=1,\forall\pi\in\mathcal{P}$).
drop-DTW (Dvornik et al., 2021) extends the set of feasible alignments, $\mathcal{P}$, to those adhering only to the monotonicity constraint. Consequently, unlike DTW, events can be dropped from the alignment process. drop-DTW finds the optimal temporal alignment while allowing for the removal of certain elements. To account for unmatched elements in the alignment cost calculation, Dvornik et al. (2021) introduce a new drop cost: $\delta\in\mathbb{R}^{+}$. [^2] In the original approach of Dvornik et al., the cost value depends on the event type. We use a constant value for the sake of simplification. The optimal matching can then be defined as follows:
$\pi^{*}=\operatorname*{argmin}_{\pi\in\overline{\mathcal{P}}}\,\langle\pi,C\rangle+\delta.\left(P_{r}(\pi)+P_{c}(\pi)\right)$
where $\overline{\mathcal{P}}\in\{0,1\}^{K}$ is the set of binary matrices satisfying only the monotonicity constraint, and $P_{c}(\pi)\in\{0,1\}^{N}$ is a vector whose $j$-th element is $1$ if $\pi_{:,j}=\bm{0}$ and 00 otherwise; $P_{r}(\pi)$ is defined similarly, but along the rows. The indices stand for rows and columns. Similar to DTW, drop-DTW can be efficiently evaluated using dynamic programming techniques.
Example 2 (Drops representation in drop-DTW).
In Example 1, the alignment of the sequences $s_{1}$ and $s_{2}$ leads to skip the $P$ event and the second $S$ event (see Figure 1). Then, the drops are represented as follows: $P_{r}=[1,1,1]$ and $P_{c}=[1,1,0,0,1]$.
@@ 3.2 Adaptation of drop-DTW for timed sequences
The proposed adaptation of drop-DTW focuses on proposing a metric between events and adding temporal constraints to the drop-DTW algorithm. The Dynamic Time Warping algorithm requires defining a cost between two events to compare sequences. To take into account both time and events, we decided to assign weights to a time distance cost $c_{t}$, and an event difference cost $c_{e}$. Having introduced the probabilistic representation of timed sequences, we define the distance between two events $e_{a}=(\bm{d}_{a}=[d^{1}_{a},\dots,d^{n}_{a}],t_{a})$ and $e_{b}=(\bm{d}_{b}=[d^{1}_{b},\dots,d^{n}_{b}],t_{b})$ by:
@@
where $p_{t},p_{e}\in\mathbb{R}_{+}$ are weights of the two terms: a distance between probabilistic events and a time gap. This metric implies that time events are in an Euclidean space of dimension $n+1$. This is of particular interest in the next section, to define average events.
We can now adapt the metric between two timed sequences. We have chosen to use drop-DTW for its ability to handle aberrant or missing events. Algorithm 1 presents the method for calculating the proposed distance between two timed sequences. It takes as input the cost matrix $C$ that computes the cost of aligning two events using the distance $d$. The algorithm finds the best alignment between sequence events to minimize the overall metric. The use of dynamic programming technique ensures the efficiency of the method. For the sake of computing efficiency, we also add constraints on the possible event pairings by allowing matches only between events that are close in time, i.e. with a delay between their occurrence lower than a predefined threshold $\tau$.
@@
$D^{+}_{0,0}\leftarrow 0$ ; $D^{+}_{i,0}\leftarrow+\infty$ ; $D^{+}_{0,j}\leftarrow+\infty$, $\forall i\in[M],\,j\in[N]$ ;
  // Match table
$D^{-}_{0,0}\leftarrow 0$ ; $D^{-}_{i,0}\leftarrow i\times\delta$ ; $D^{-}_{0,j}\leftarrow j\times\delta$, $\forall i\in[M],\,j\in[N]$ ;
  // Drop table
$D_{0,0}\leftarrow 0$; $D_{i,0}\leftarrow D^{-}_{i,0}$; $D_{0,j}\leftarrow D^{-}_{0,j}$, $\forall i\in[M],\,j\in[N]$;
  // Optimal solution table
@@
4
5            else
                   $D_{i,j}\leftarrow+\infty$ ;
                    // Unsatisfied temporal constraints $\Rightarrow$ impossible match
@@
Considering the sequences from Example 1, sequences $s_{1}$ and $s_{3}$ are very similar, but sequences $s_{1}$ and $s_{2}$ does not seem so different if we ignore some of the events. In the proposed alignment between $s_{1}$ and $s_{2}$, the drop-cost was lower than the cost of pairing $P$ in the sequence $s_{2}$ with another event of $s_{1}$. Thus, it was preferable to ignore this event. This is directly related to the definition of the distance between a physiologist session and other types of event. In the case of the second surgery event in $s_{2}$, it was too far in time from the one in $s_{1}$ according to the $\tau$ constraint. Then, against, it was preferable to ignore it in the final alignment.
@@ 4 Construction of an Average Timed Sequence
Computing an average timed sequence consistent with the notion of similarity is essential for clustering algorithms. In the following, we propose a method inspired by the DBA algorithm (Petitjean and Gançarski, 2012) to create average sequences in terms of drop-DTW, taking into account the possibility of removal. DBA addresses the problem of generating an average time series in terms of DTW – which does not hold distance properties. In this case, the generalized average is defined as a time series that minimizes the global inertia (see Equation 1).
DBA iteratively refines a (randomly initialized) series to minimize its quadratic distance (DTW) with respect to the series to be averaged. At each iteration, the algorithm performs “vertical” centroids of the aligned elements of the time series. The definition of a “vertical” centroid corresponds to averaging series real values. Our approach enables the adaptation to timed sequences thanks to the representation of events as distributions of types.